# Recursive Training Loops in LLMs: How training data properties modulate distribution shift in generated data?

### v4

@@ B.2 Dataset details
Throughout the study we use the following five datasets. The 100M_tweets  [^2] https://huggingface.co/datasets/enryu43/twitter100m_tweets ($C$-BY-4.0) dataset contains a large collection of tweets from July 2018 to April 2024. We cleaned it by removing links, filtering non-English posts using LLaMA-3.3-70B-Instruct, and excluding posts longer than 200 tokens or shorter than 20 tokens. Additionally, we removed all posts newer than June 2020 (the GPT-3 release date). The final cleaned dataset consists of 2 million posts. The senator_tweets  [^3] https://huggingface.co/datasets/m-newhauser/senator-tweets dataset contains all tweets made by United States senators during the first year of the Biden Administration (2021). We cleaned it by removing links and posts shorter than 10 tokens. The final cleaned dataset consists of 94878 posts. The wikipedia (Foundation, )  [^4] https://huggingface.co/datasets/wikimedia/wikipedia ($C$-BY-SA-3.0) dataset was created by compiling and cleaning articles from Wikipedia dumps  [^5] https://dumps.wikimedia.org/ in November 2023. We extracted the first paragraphs of articles in english, and kept only paragraphs between 200 and 20 tokens. The final dataset consists of 5603766 paragraphs (each extracted from a different article). The reddit_submissions   [^6] https://huggingface.co/datasets/HuggingFaceGECLM
@@
(arXiv.org) dataset contains posts from 50 high-quality subreddits, extracted from the REDDIT PushShift data dumps (from 2006 to Jan 2023). We pre-processed this dataset by merging post titles with bodies, sampling 25000 posts from each Subreddit, removing those that have [deleted] or [removed] tags, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1243794 posts. The webis_reddit  Völske et al. (2017) ($C$-BY) dataset contains preprocessed posts from the Reddit dataset (Webis-TLDR-17). We pre-processed this dataset by merging titles with bodies, removing "tldr" tags, removing posts that are marked as "nsfw" or "+18", removing duplicates, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1458003 posts.
@@ B.3 LLM-as-a-judge validation
To confirm the correctness of our judge, we compare it to the judge from Chen et al. (2023), which was shown to outperform many other quality metrics. We use the same experimental setup as in Chen et al. (2023) and compare to numbers extracted from Table 10 in Chen et al. (2023). Table 2 shows the correlations for the two judges with human annotations of the OpenMEVA-ROC dataset Guan et al. (2021) of generated stories. We can see that our judge matches that of Chen et al. (2023) justifying it’s use in our experiments.
@@ B.3.1 Data properties selection
  • Semantic Diversity: We compute several semantic diversity metrics as pairwise cosine diversity in the embeddings of the stella_en_1.5B_v5 model (Zhang et al., 2024a). Cosine diversity computes the pairwise diversity between all data points, and k-$n$ Cosine diversity computes the pairwise diversity for the nearest n neighbors (we use this metric with n=50 and n=1000).
@@ B.4 Clustering
To obtain data to be used for regression analysis, we create a number of subsets from each of the datasets. This is done by the following procedure. A 2D UMAP is fit on the 90k text embedded with stella_en_1.5B_v5 model (Zhang et al., 2024a). A series of clustering methods (dbscan, hdbscan, gmm, k-means) with different hyperparameters are done separately, each annotating the 90k samples. Then the rest of the dataset is annotated by k-$n$, with k=1. This is done in two ways, with and without excluding the noise cluster for the k-$n$ classifier. This gives us a total of 120 different clusterings of the dataset. Then for each clustering 10 clusters are taken, if there are not enough clusters over the size of 60k. The remaining clusters are constructed by merging smaller clusters, either by uniformly sampling which cluster to merge or by iteratively merging the cluster that is the furthest away from the currently merged cluster. As this results in a large number of clusters (e.g. 1088 for webis_reddit ), we then obtain the final cluster set by subsampling 200 clusters.
@@ C.2.1 Effect on quality
Figure 9 shows the quality values of iterative chains for different quality levels. As in section 4.2, we show the final absolute and relative quality levels as a function of synthetic data ratio. Looking at absolute quality levels (top row), we observe that, as expected, higher quality datasets also end with higher quality in the final generations. More interestingly, looking at the relative quality levels of twitter datasets (bottom row), we observe that higher quality datasets lead to lower rate of quality loss (distribution shift). That is, not only does higher input data quality increase the quality of the final generated dataset, it also decreases the percentage of original quality lost due to recursive training. Furthermore, focusing on lower ratios (<1/4<1/4) of the 100M_tweets dataset, we observe that the higher quality dataset are more robust to increasing synthetic data ratio. For the quality of 8080, major quality losses are observed only at $r=1/4$, while for lower qualities it is observed already at $r=1/8$. Finally, it should also be noted that the 100M_tweets dataset with quality 2020 does not appear to lead to significant shifts with higher synthetic data ratios. Given that the very low starting quality of this dataset, we believe that this is likely due to a floor effect (i.e. there are no losses in quality because the initial dataset was already close to the lower bound. Curiously, on both Reddit datasets, we do not observe strong differences when manipulating quality, implying that there are other factors than quality influencing collapse. A similar study focusing on losses of diversity is presented in Appendix C.2.2, where an effect is observed only on the senator_tweets dataset. Overall, this experiment shows that, in some conditions, high-quality datasets lead to increased robustness to distribution shifts.
@@ C.2.2 Effect on diversity
C.4 Additional experiments on the distribution shift of political lean

### v5

@@ B.2 Dataset details
Throughout the study we use the following five datasets. The 100M_tweets  [^2] https://huggingface.co/datasets/enryu43/twitter
100m_tweets
($C$-BY-4.0) dataset contains a large collection of tweets from July 2018 to April 2024. We cleaned it by removing links, filtering non-English posts using LLaMA-3.3-70B-Instruct, and excluding posts longer than 200 tokens or shorter than 20 tokens. Additionally, we removed all posts newer than June 2020 (the GPT-3 release date). The final cleaned dataset consists of 2 million posts. The senator_tweets  [^3] https://huggingface.co/datasets/m-newhauser/senator-tweets dataset contains all tweets made by United States senators during the first year of the Biden Administration (2021). We cleaned it by removing links and posts shorter than 10 tokens. The final cleaned dataset consists of 94878 posts. The wikipedia (Foundation, )  [^4] https://huggingface.co/datasets/wikimedia/wikipedia (CC-BY-SA-3.0) dataset was created by compiling and cleaning articles from Wikipedia dumps  [^5] https://dumps.wikimedia.org/ in November 2023. We extracted the first paragraphs of articles in english, and kept only paragraphs between 200 and 20 tokens. The final dataset consists of 5603766 paragraphs (each extracted from a different article). The reddit_submissions  [^6] https://huggingface.co/datasets/HuggingFaceGECLM
@@
(arXiv.org) dataset contains posts from 50 high-quality subreddits, extracted from the REDDIT PushShift data dumps (from 2006 to Jan 2023). We pre-processed this dataset by merging post titles with bodies, sampling 25000 posts from each Subreddit, removing those that have [deleted] or [removed] tags, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1243794 posts. The webis_reddit Völske et al. (2017) (CC-BY) dataset contains preprocessed posts from the Reddit dataset (Webis-TLDR-17). We pre-processed this dataset by merging titles with bodies, removing "tldr" tags, removing posts that are marked as "nsfw" or "+18", removing duplicates, and removing posts longer than 200 tokens or shorter than 20 tokens. The final cleaned dataset consists of 1458003 posts.
@@ B.3 LLM-as-a-judge validation
To confirm the correctness of our judge, we compare it to the judge from Chen et al. (2023), which was shown to outperform many other quality metrics. We use the same experimental setup as in Chen et al. (2023) and compare to numbers extracted from Table 10 in Chen et al. (2023). Table 2 shows the correlations for the two judges with human annotations of the OpenMEVA-ROC dataset Guan et al. (2021) of generated stories. We can see that our judge matches that of Chen et al. (2023) justifying its use in our experiments.
@@ B.3.1 Data properties selection
  • Semantic Diversity: We compute several semantic diversity metrics as pairwise cosine diversity in the embeddings of the stella_en_1.5B_v5 model (Zhang et al., 2024a). Cosine diversity computes the pairwise diversity between all data points, and k-nn Cosine diversity computes the pairwise diversity for the nearest n neighbors (we use this metric with n=50 and n=1000).
@@ B.4 Clustering
To obtain data to be used for regression analysis, we create a number of subsets from each of the datasets. This is done by the following procedure. A 2D UMAP is fit on the 90k text embedded with stella_en_1.5B_v5 model (Zhang et al., 2024a). A series of clustering methods (dbscan, hdbscan, gmm, k-means) with different hyperparameters are done separately, each annotating the 90k samples. Then the rest of the dataset is annotated by k-nn, with k=1. This is done in two ways, with and without excluding the noise cluster for the k-nn classifier. This gives us a total of 120 different clusterings of the dataset. Then for each clustering 10 clusters are taken, if there are not enough clusters over the size of 60k. The remaining clusters are constructed by merging smaller clusters, either by uniformly sampling which cluster to merge or by iteratively merging the cluster that is the furthest away from the currently merged cluster. As this results in a large number of clusters (e.g. 1088 for webis_reddit ), we then obtain the final cluster set by subsampling 200 clusters.
To clarify, from the 200 clusters created for each dataset, some are created in a straight forward way (e.g. with k-nn) which will create the nice narrow clusters as the reviewer describes. However, as discussed in this section, some clusters are created in a slightly more complex way, precisely to address this issue. Some of the used clustering methods create very small clusters (miniclusters) which are then merged to create bigger final clusters to be used in the experiments. This is sometimes done by randomly selecting the miniclusters to merge, and sometimes by iteratively adding the farthest away minicluster from the already merged miniclusters. Therefore, the final set of 200 clusters will be quite diverse.
@@ B.5 Fine-tuning procedure
B.6 Iterative chain pseudocode
Figure 8 shows the pseudocode describing the iterative chain experiment used. Each iteration a fresh base model is selected and finetuned.
1function iterative_chain(human_data, ai_ratio):
2 base_models = [llama-3.2-1B, Qwen2.5-1.5B, SmolLM-1.7B, Falcon3-1B-Base]
3 accumulated_data_pool = []
4 for i in 20:
5 iteration_base_model = sample(pretrained_models)
6
7 if i == 0:
8 training_set = sample(human_data, 8000)
9 else:
10 training_set = sample(accumulated_data_pool, 4000)
11
12 iteration_ft_model = iteration_base_model.train_with_lora(training_set)
13
14 iteration_new_data = iteration_ft_model.generate(4000*ai_ratio)
15 iteration_human_data = human_dataset.sample(4000*(1-ai_ratio))
16
17 accumulated_data_pool.add(iteration_new_data)
18 accumulated_data_pool.add(iteration_human_data)
Figure 8: Iterative chain pseudocode
B.7 Discussion on corrections for multiple comparisons
We report uncorrected p-values in our regression analyses as our goal is not to test specific hypotheses about individual coefficients, but rather to identify general patterns of robust influence across datasets and conditions. On a related note, variable selection was conducted prior to these analyses and applied to newly collected data (last sentence in Appendix B.3.1). We therefore believe that the risk of multiple testing bias is minimal.
In other words, we address the issue of Type I errors in the way we interpret our results and make conclusions (i.e. in a way that is robust to potential Type I errors): the fact that we consistently find the same predictors to be significant across different experiments gives a high confidence that those are not statistical artifacts.
@@ C.2.1 Effect on quality
Figure 10 shows the quality values of iterative chains for different quality levels. As in section 4.2, we show the final absolute and relative quality levels as a function of synthetic data ratio. Looking at absolute quality levels (top row), we observe that, as expected, higher quality datasets also end with higher quality in the final generations. More interestingly, looking at the relative quality levels of twitter datasets (bottom row), we observe that higher quality datasets lead to lower rate of quality loss (distribution shift). That is, not only does higher input data quality increase the quality of the final generated dataset, it also decreases the percentage of original quality lost due to recursive training. Furthermore, focusing on lower ratios (<1/4<1/4) of the 100M_tweets dataset, we observe that the higher quality dataset are more robust to increasing synthetic data ratio. For the quality of $80$, major quality losses are observed only at $r=1/4$, while for lower qualities it is observed already at $r=1/8$. Finally, it should also be noted that the 100M_tweets dataset with quality $20$ does not appear to lead to significant shifts with higher synthetic data ratios. Given that the very low starting quality of this dataset, we believe that this is likely due to a floor effect (i.e. there are no losses in quality because the initial dataset was already close to the lower bound. Curiously, on both Reddit datasets, we do not observe strong differences when manipulating quality, implying that there are other factors than quality influencing collapse. A similar study focusing on losses of diversity is presented in Appendix C.2.2, where an effect is observed only on the senator_tweets dataset. Overall, this experiment shows that, in some conditions, high-quality datasets lead to increased robustness to distribution shifts.
@@ C.2.2 Effect on diversity
C.4 Regression models’ explained variances
Here we provides $r^{2}$ measures for the linear regression models used in experiments in sections 4.3 and 4.4. Table 3 corresponds to table 1 from section 4.3. Table 4 corresponds to Figure 4 from section 4.4. We can see that the considered properties can partially account for observed effect, many more effects remain to be uncovered (in particular in the 100M tweets and wikipedia datasets).
$\mathbf{r^{2}}$ All Webis 100M Tweets Reddit Submissions Wikipedia
Synthetic data ratio 1/8 1/4 1/8 1/4 1/8 1/4 1/8 1/4
Semantic Diversity 0.293 0.244 0.541 0.043 0.188 0.109 0.502 0.147 0.105
Quality 0.737 0.482 0.353 0.193 0.324 0.143 0.564 0.249 0.276
Table 3: $R^{2}$ results corresponding to Table 1.
Domain $R^{2}$ (Diversity) $R^{2}$ (Quality)
Webis 0.340 0.591
100M Tweets 0.532 0.611
Wikipedia 0.241 0.363
Table 4: $R^{2}$ results corresponding to Figure 4.
C.5 Additional experiments on the distribution shift of political lean